UT 20131120 - 20131212 LUCI2 Closed-dome Commissioning

Observer: LUCI-team
Telescope Operator: various
LBTO Support:

Details

November 18-21 2013

Exposure times covering all data obtained for Com1:

With the faster readout times:
  • 1.370 1 x blind blind
  • 1.651 10 x K
  • 2.000 1 x No Filter (clear clear)

With the default readout time:
  • 2.740 1 x HK_spec
  • 2.750 1 x K
  • 3.000 30 x zJ_spec
  • 4.000 1 x Ks
  • 5.000 12 x J
  • 8.000 1 x Ks
  • 10.000 5 x Ks
  • 12.000 1 x z
  • 15.000 1 x HK_spec
  • 20.000 5 x OH_1060
  • 30.000 3 x zJ_spec
  • 40.000 1 x J
  • 60.000 3 x HK_spec
  • 300.000 1 x blind blind

Created Darks_L2com1.xml script to obtain the darks listed above with the default readout times. Filters and NCOADDS do not matter for darks, so a reasonable number was selected for each DIT. Check NCOADDS in header before using these data to dark-subtract any of the Com1 data!

DThompson tried to execute the script Darks_L2com1 but repeatedly failed.
Stopped at first item
JHeidt tried it later on again. Script stopped repeatedly at item14 out of 15.
After restarting GEIRS, GEIRS services and readout control script went well
just for items 14 and 15. Complete set for script Darks_L2com1_mod.xml
(used initial setup by Dave but modified to DIT=3 in item 1 and 14 for LIR
and DIT=7 in item 15 for MER. Files are #0092-0165 in folder 20131121.
Why this is so is unclear. APramskiy is working on it.

November 22 2013

DThompson managed to run most of the Darks_L2com1.xml script through completely. Full start up following notes supplied by JHeidt and APramskiy worked. Script ran through to item #14 out of 15 and then stopped. Message GUI info includes:
  • GEIRS received response with error (2013-11-22 17:28:51)
  • Readout error in GEIRS application (2013-11-22 17:28:51)
  • Save FITS file error in GEIRS application (2013-11-22 17:28:53)
  • Error committing setup to readout manager (2013-11-22 17:28:53)

Majority of the dark data was taken repeatedly as we tried to get the script to run through completely. The last set was luci2.20131122.NNNN.fits (NNNN > 300).

UT 20131123 - Calibration Unit flatfields for the W3 data
  • OB: ~/COMMISSIONING_STUFF_ONLY/CUFlats_JHK.xml
  • luci2.20131122.NNNN

0073 1x6.0s blind filter to test
0074 - 0078 1x6.0s J filter ~10k ADU
0079 - 0083 1x4.0s H filter ~17k ADU
GEIRS error when taking the K data (no data taken)
- stopped and restarted GEIRS Readout, Readout manager, and GEIRS itself (because I could not read out the array from the GEIRS GUI

Script modified to take lamp-off and lamp-on data: CUFlats_JHK_v2.xml 0084 0093 1x6.0s J lamp-off below 100 ADU
0094 0103 1x6.0s J lamp-on ~10k ADU
Error commiting readout setup at item 3
- stopped and restarted GEIRS Readout, Readout manager, and GEIRS itself (because I could not read out the array from the GEIRS GUI
skip 2 items, commit 2 items
0104 0113 1x3.0s H lamp-off below 100 ADU
0114 0123 1x3.0s H lamp-on ~8k ADU
0124 0133 1x3.0s K lamp-off ~14k ADU
0134 0143 1x3.0s K lamp-on ~24k ADU

Since that set finished, reset queue and commit all 6 items.
0144 0153 1x6.0s J lamp-off below 100 ADU
0154 0163 1x6.0s J lamp-on ~8k ADU
0164 0173 1x3.0s H lamp-off ~14k ADU
0174 0183 1x3.0s H lamp-on ~24k ADU
0184 0193 1x3.0s K lamp-off ~14k ADU
0194 0203 1x3.0s K lamp-on ~24k ADU
Script completed!
Reconfigure: blind mask and filters, N3.75, mirror

Notes:
  • Lamp shows as ON in the UI, not on READ2 GUI (works if you refresh)
  • Completion of the item does not turn off lamp
  • No lamp-off data taken with the lamp-on data automatically

November 29 2013

LBTB-nigth weathered out, started with grating calibration of G150, finished it, but with several problems
Got frequently "error comitting setup to instrument manager error",
mostly by "unable to initialize compensation mirror", stop/restart SW didn't help
In the end FW2 got stuck

November 30 2013 Again LBTB-nigth weathered out. Alexander put in a new SW-version.
Everything went smoothly without any single problem.
All gratings/order are well calibrated now and ready for COM#2

G150:
Lambda/Volt/Lamp
1.9823 +1.010 Ar
1.9885 +0.885
2.0323 +0.580
2.0622 +0.310
2.0992 -0.037
2.1339 -0.372
2.1540 -0.572
2.1714 -0.741 Ne
2.2083 -1.100 Ar
2.2537 -1.536 Ne
2.3140 -2.099 Ar
G200zJ:
0.9227 0.420 Ar
0.9660 1.070
0.9787 1.258
0.9923 1.460 Xe
1.0124 1.730
1.0472 2.270 Ar
1.0676 2.560
1.0841 2.795 Xe
1.1088 3.145
1.1394 3.567 Ne
1.1516 3.700 Ar
1.1684 3.946
1.1988 4.353 Ne
G200HK:
1.5049 -1.85 Ar
1.5292 -1.693
1.5994 -1.258
1.6524 -0.916
1.6945 -0.623
1.7437 -0.277
1.7904 0.0550
1.8414 0.415
1.8777 0.685 Xe
1.9823 1.437 Ar
2.0323 1.800
2.0622 2.020
2.0992 2.294
2.1339 2.542
2.1540 2.688
2.1714 2.815 Ne
2.2083 3.077 Ar
2.2537 3.388 Ne
2.3138 3.800 Ar
G210z
0.8427 4.014 Ar
0.8471 3.821 Ne
0.8569 3.571 Ar
0.8822 2.716 Xe
0.8955 2.270
0.9048 1.947
0.9125 1.673 Ar
0.9227 1.300
0.9660 -0.368
0.9787 -0.858
0.9923 -1.384 Xe
1.0032 -1.763
1.0124 -2.070
G210J:
1.1516 1.425 Ar
1.1684 0.885
1.1770 0.585 Ne
1.2116 -0.490 Ar
1.2143 -0.578
1.2347 -1.204
1.2406 -1.386
1.2491 -1.637
1.2706 -2.270
1.2806 -2.555
1.2937 -2.915
1.3012 -3.114
1.3242 -3.683
G210H:
1.5085 2.014 Ar
1.5199 1.739
1.5251 1.617 Ne
1.5346 1.398 Ar
1.5425 1.200 Xe
1.5904 0.096 Ar
1.5994 -0.118
1.6058 -0.264 Xe
1.6184 -0.560 Ar
1.6441 -1.150
1.6524 -1.347
1.6745 -1.840
1.6945 -2.281
1.7119 -2.745 Ne
1.7330 -3.080 Xe
1.7475 -3.310 Ar
G210K:
2.0622 4.329 Ar
2.0992 3.965
2.1047 3.905 Ne
2.1339 3.589 Ar
2.1540 3.360
2.1714 3.145 Ne
2.2083 2.686 Ar
2.2580 2.067 Ne
2.3096 1.197 Ar
2.3519 0.830 Ne
2.3827 0.100 Ar

December 1 2013

Weather at LBT is now bad, will use time for internal flats.
Plan is to do all filters with N3.75-camera and BB-filters
with N1.8 camera. Always 10 lamp on/off images.

LT: 12:45 Startup of system => OK
Files are in dir 20131130, luci2.20131130.*
Files 1-14 tests for focus-script
Start 13:15
N3.75
K: 3sec, lamp3, on=15-24, off=25-34, 24k/14k
Ks: 4sec, lamp3, on=35-44, off=45-54, 18k/7k
H: 5sec, lamp3, on=55-64, off=65-74, 10k/0k
J: 6sec, lamp3, on=75-84, off=85-94, 6k/0k
z: 2.75sec, lamp2, 95-104 crap, on=105-114, off=115-124, 30k/0k
H2: 2.75sec, lamp2, on=125-134, off=135-144, 19k/10k
BR_gam: 3sec, lamp2, on=145-154, off=155-164, 20k/0.5k
OH_1060: 10sec, lamp2, on=165-174, off=175-184, 12k/0k

=> GEIRS error, when starting 10 exp von ReadOut Manager
(save FITS-error in GEIRS application)
=> Stop/Start GEIRS, GEIRS service, ReadoutManager => OK

Y1: 3sec, lamp2, on=185-194, off=195-204, 17k/0k
Y2: 2.75sec, lamp2, on=205-214, off=215-224, 27k/0k
P_gam: 3sec, lamp2, on=225-244 (erroneously twice), off=245-254, 5k/0k
P_beta: 5sec, lamp2, on=255-264, off=265-274, 15k/0k
OH_1190: 7sec, lamp2, on=275-284, off=285-294, 15k/0k
FeII: 2.75sec, lamp2, on=295-304, off=305-314, 19k/0k
J_low: 20sec, lamp3, on=315-324, off=325-334, 11k/0k

Images 335/336 crap

J_high: 20sec, lamp3, on=337-346, off=347-356, 12k/0k
HeI: 7sec, lamp2, on=357-366, off=367-376, 15k/0k

N1.8:
z: 12sec, lamp3, on=377-386, off=387-396, 10k/0k
J: 2.75sec, lamp3, on=397-406, off=407-416, 18k/0k
H: 2.75sec, lamp3, on=417-426, off=427-436, 30k/0k
Ks: 2.75sec, lamp3, on=437-446, off=447-456, 43k/0k (probably too bright)
K: 2.75sec, saturated, no chance

December 5 2013

Weather is bad at LBT, can do some tests. Dave tried to run a
Dark-script but failed again. Took over and do some grating-repro tests.
Lights on dome are off, scope in zenith
Connection to mountain is veeeeery slow
LT 8:10, start-up system, seems OK
Files are luci2.20131204.****

ROTANGLE=210

8:45 Grat_tilt_repro_G210, #14-19
8:47 Grat_tilt_repro_G200, #20-25
8:49 Grat_tilt_repro_G150, #26-27, grating did not stabilize
at obsitem #3, 2.17mue (same tilt as file #25)
8:52 Redo Grat_tilt_repro_G150, did not stabilize again
this time at obsitem #4, 2.22mue, files #28-30
8:56 Redo Grat_tilt_repro_G150, failed at obsitem #1
8:58 Redo Grat_tilt_repro_G150, failed at obsitem #1
9:00 Redo Grat_tilt_repro_G150, failed at obsitem #3
9:02 Redo Grat_tilt_repro_G150, failed at obsitem #2
Doesn't make sense, stopped
9:04 Grat_select_repro, files #35-37 again G150 didn't want to stabilize
9:06 Redo Grat_select_repro, same problem again. #38-41
=> Needs to be checked, stabilizes sometimes but mostly not at
2.17mue (but also at 2.12/2.22mue.
=> See whether this problem holds for other ROTANGLES as well.

ROTANGLE=300

9:17 Grat_tilt_repro_G210, #42-47
9:21 Grat_tilt_repro_G200, #48-53
9:25 Grat_tilt_repro_G150, #54-27, grating did not stabilize at 2.17mue
9:28+ redo several times, always stopped at some point
9:32 Grat_select_repro, just to double-check, files #57-63 same as before

ROTANGLE=120

10:04 Grat_tilt_repro_G210, stopped, issue with FW2 when moving to K-filter
10:04 Grat_tilt_repro_G210, #64-
10:08 Issue with FW2 when moving to K-filter as yesterday, need to
wait for Walter to fix it, back to ROT=210, blind mask in FPU, CU out

December 6 2013

Weather is again bad, do tests.
LT 08:50 Startup system, everything seems to work now.
Do first check of G150 at ROTANGLE=120,30
Files are luci2.20131205.****

ROTANGLE=120

9:05 Grat_tilt_repro_G150, got immediately error
when moving FW2 to pos 7 (Ks) => need to wait for Walter
=> is fixed now, try again:
10:17 Grat_tilt_repro_G150, need to initialize instr.manager,
can not do it, problems in ini compensation mirror?
Didn"t have this problem yesterday!!!
logout/login and retry, didn"t help.
Called Alexander: problem is apparently telescope service,
which had to be restartet as well as DD, comp.mirror, instr.manager & scheduler
11:00 Grat_tilt_repro_G150, stops already at first obsitem,
...can not stabilize grating
11:15 Try to do start Grat_tilt_repro_G200 => FW2 issue,
try to fix that by my own, OK, successfull
11:29 restart Grat_tilt_repro_G200, files #3-9, OK
11:33 Grat_tilt_repro_G150, do not stabilize...

ROTANGLE=30

11:37 Grat_tilt_repro_G210, #10-15
11:45 Grat_select_repro, G150 can not stabilize again
==> No stabilization of G150 at any ROTANGLE

Back to ROTANGLE=210

Focus-determination:
N180: stepsize = 5, number of steps = 11
K, start = -325, no lamp, 3sec, #18-28
N3.75: stepsize = 10, number of steps = 11
K, start = -420, no lamp, 3sec, #29-39
Ks, start = -420, no lamp, 5sec, #40-45
ReadoutManager got stuck, redo,
ini Inst.Manager, FW2 got stuck, recover,
Ks, start = -420, no lamp, 5sec, #47-57
H, start = -420, lamp3, 3sec, #58-68
J, start = -410, lamp3, 3sec, #69-79
z, start = -400, lamp2, 3sec, #80-81
ROM got stuck again,
z, start = -400, lamp2, 3sec, #84-94
OH_1190, start = -390, lamp2, 3sec, #95-105
FeII, start = -420, lamp2, 5sec, #106-116
Y1, start = -390, lamp2, 5sec, #117-127
J_low, start = -390, lamp2, 5sec, #128-138
OK, done

Try to execute Daves script DN:
14:05 Script started, files #139-194.
Crazy, took all requested images, but the last 16 were not with long exp.times
as requested but only with minimum DIT.
Restart the last 4 obsitems (texp=60,300,600,900sec).
Same again, cyctle-time always 13sec!
Files #195-210
=> Alexander need to check

Files #211-213, Tests with G150, will not stabilize properly,
increased the time to let it stabilize to 60sec instead
of 30sec did not help much.
Walter/Michael will have a look on this

December 10 2013

Attempt to re-run Dark_RN.xml script

GEIRS gui startup normal, image taken
LMC startup normal. Readnout Manager not running
Stop GEIRS Readout service
Stop Scheduler
Start GEIRS Readout service
Start Readout Manager
Start Scheduler

Start Observer Execution Panel
LUCI2 gui comes up
READ2 gui comes up
Read of detector normal

Files: luci2.20131209.NNNN.fits

0001 Test exposure from READ2 gui

Load Dark_RN.xml
commit all 6 items (UT 18:13)

0002 0021 1x2.7399s LIR
0022 0041 1x6.8515s MER
0042 0045 1x60s MER
0046 0049 1x300s MER
0050 0053 1x600s MER
0054 0057 1x900s MER

Script hung at 0051 with a GEIRS error
Jochen restarted things

Rollover to next night's naming:
luci2.20131210.NNNN.fits

0001 ?? test exposure?
May be a third 600s exposure started before noon and running past the rollover

Reset Dark_RN.xml queue
skip 4 items
commit (last) 2 items

0002 0005 1x600s
0006 0009 1x900s

21:50 Queue reset and full 6 items committed

0010 0029 1x2.7399
0030 0049 1x6.8515
0050 0053 1x60s
0054 0057 1x300s
0058 0061 1x600s -> script hung after 0058 again
0062 0065 1x900s (not taken)

01:23 Redo Dave's Dark_RN-Script to see if it again hungs.

0059 0078 1x2.7399
0079 0098 1x6.8515
0099 0102 1x60s
0103 0106 1x300s
0107 0110 1x600s
0111 0114 1x900s

Went through without problems.

Focus determination. N1.8 camera, K-filter, 3sec, Optik sieve,
start = -325, steps = 5, number of steps 22 => 0115 - 0136

Image with "persistence diagonal mask", 0137

Image with N1.8 + 1.0" ACF-slit to determine tilt of slits, 0138.
Tilt dx is 1pix (0.12") for 85pix (10.2") in y

December 12 2012

Startup OK. Both READ2 gui and GEIRS read out OK

Files: luci2.20131211.NNNN.fits

0001 0020 1x2.7399s LIR for readnoise (temporal and spatial)
0021 0040 1x6.8515s MER for readnoise (temporal and spatial)

First two items in Dark_RN.xml completed without problems.

Configure with pointing mask...OK
Highbay flood lights off in enclosure
Configure Br_gam + K filters

Take test image through pointing mask...ERROR
GEIRS received response with error 255

Took manual exposure with GEIRS gui...OK
Took manual image with READ2 gui...OK
seems OK

Humidity rose at sunset and we had to close.
01:14 Still closed for humidity
Opened, clearing

JHill and DMiller testing some stuff from Tucson

### Transform calibration data...

# On sky but with a blind filter in (use as dark to see hole circle)
0041 0044 4 x 3.0s LIR saved "normal"
0045 4 x 3.0s LIR saved "integrated"

See JHill's log on wiki...config problem on AGw2

Testing execution of the Dark_RN.xml script
Execution panel hung, various partial restarts did not help. Had to start as
though it was a new login.

Dark_RN.xml (commit all 6 items). Note that instrument will reconfigure from
the pointing mask and inset the calib unit.

UT08:33 0046...0102 Script completed all elements

Calib unit moved out of beam

-- JochenHeidt - 13 Dec 2013
Topic revision: r9 - 13 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback